x=y,2x + 2y =120

Please help mee
Thanks in advance!:)

Answer by jim_thompson5910(35256)

2x + 2y =120

2y + 2y =120 ... plug in x = y (ie replace x with y)

4y =120

y =120/4

y =30

So x = 30 and y = 30